What is the value of each of the angles of a triangle whose sides are 148, 190 and 94 cm in length?
Answer #1

here,

sides , a = 1.48 m , b = 1.9 m , c = 0.94 m

let the angles be A , B and C

so,

cos(C) = (a^2 + b^2 - c^2) /( 2 * a * b)

cos(C) = ( 1.48^2 + 1.9^2 - 0.94^2) / ( 2 * 1.48 * 1.9)

C = 29.04 degree

and

cos(B) = ( a^2 + c^2 - b^2)/( 2 * a * c)

cos(B) = ( 1.48^2 + 0.94^2 - 1.9^2) / ( 2 * 1.48 * 0.94)

B = 101.11 degree

and

A + B + C = 180 degree

A + 29.04 + 101.11 = 180

A = 49.85 degree